OpenStudy (anonymous):

Hey looking for help ASAP have a test really soon, cant understand this problem. Will post below, calculus related (integrals). Medal ofcourse for answer, thanks a lot and if it can be answered fast thatd be great.

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin^3(2x)\cos^8(2x)dx\]

OpenStudy (anonymous):

I have tried so many ways, using a u sub of cos(2x) and changing sin(2x) to 2sinxcosx and even sin^2to 1/2(1-cos(2x))

OpenStudy (anonymous):

Anyone?

OpenStudy (tkhunny):

What Parts have you tried? Doesn't seem TOO tricky (albeit rather annoying). \(\int\sin^{3}(2x)\cos^{8}(2x)\;dx = \int\sin^{2}(2x)\;d\left[(-1/18)\cos^{9}(2x)\right]\) \(\;\;= -(1/18)\cos^{9}(2x)\sin^{2}(2x) + \int\left[(2/9)\cos^{10}(2x)\right]\cdot \sin(2x)\;dx\) This was quite fruitful, though it may not appear so. Integration by Parts a second time and you will have it.

OpenStudy (tkhunny):

Of course, it may be WAY easier to trade to since for two cosines and make your life easier. \(\sin^{3}(2x)\cos^{8}(2x) = \sin(2x)\left(1-\cos^{2}(2x)\right)\cos^{8}(2x)\) \(\;\; = \sin(2x)\left(\cos^{8}(2x) - \cos^{10}(2x)\right)\) Ah, yes. Now the world is a better place.

OpenStudy (anonymous):

@tkhunny thanks but the answer the book gives is 1/396 cos^9(2 x) (-13+9 cos(4 x)) is this equivelent?

OpenStudy (anonymous):

and yea i have got it to that form before, but then coildnt proceed

OpenStudy (anonymous):

@tkhunny how does it help find the integral?

OpenStudy (anonymous):

still trying to find it anyone can help?

OpenStudy (tkhunny):

That last form? Try \(u = \cos(2x)\). Gotta keep your eyes open...

OpenStudy (anonymous):

i have that tried that multiple times i dont think thats gives the right answer...

OpenStudy (tkhunny):

?? That makes no sense at all. If you do it right, you will get the right answer. Try not to forget that an indefinite integral has that pesky "+C" on the end. It is NOT optional. It is what makes \(\sin^{2}(x) + C\;and\;\cos^{2}(x) + C\) the SAME result!

OpenStudy (anonymous):

Okay so should it be -1/2 u^9/9-u^10/10+c and subing in cos(2x) as u?

OpenStudy (anonymous):

u^11/11**

OpenStudy (tkhunny):

And, yes, the book's answer is equivalent. Why not just do it correctly, so you can follow your work, and forget about what the book says. So far, I have three equivalent expressions. Shall we start a collection?

OpenStudy (anonymous):

Thank you, that is so nice to know

OpenStudy (anonymous):

Sure what are the other forms? I want to see if I had got them as well while working on this problem

OpenStudy (tkhunny):

The lead directly from the Parts Twice and the Trig + Substitution. You may explore it. Obviously, with that \(\cos(4x)\), in there, the book used a different trig identity somewhere along the line.

OpenStudy (anonymous):

". . . just do it correctly, . . ., and forget about what the book says. " I think I need to frame this quote and hang it on my wall.